# Forecasting Influenza Hospitalizations Using a Bayesian Hierarchical Nonlinear Model with Discrepancy

### v1

@@ 3.1 ILI Model
The proposed model for ILI for any location is given in (1). Here $ILI_{s,w}$ is the ILI for flu season $s$ and week $w=1,2,...,W$, where $W=52$ or $W=53$, depending on how many Sundays there are in a given season. The ILI is a proportion, so the Beta random variable is a natural selection for modeling. Under the parameterization in of the Beta distribution used in (1) the expected value is $\pi_{s,w}$ and the variance is $\pi_{s,w}(1-\pi_{s,w})/(1+\kappa_{s})$, making $\kappa_{s}$ a scale parameter. The nonlinear function $f_{\theta_{s}}(w)$ captures the trajectory of the ILI, and $\gamma_{w}$ is a discrepancy term for capturing the systematic patterns which $f_{\theta_{s}}(w)$ does not capture.
@@
$\text{logit}(\pi_{s,w})$ $=f_{\theta_{s}}(w)+\gamma_{w}$
In Osthus et al. Osthus et al. (2019) $f_{\theta_{s}}(w)=\text{logit}(I_{s,w})$ where $I_{s,w}$ is the infectious compartment of the SIR model from (2) in section 3.2. In Ulloa Ulloa (2019) $f_{\theta_{s}}(w)=ASG_{\theta}(w)$ from (3) in section 3.3. In Ulloa Ulloa (2019) modeling is done hierarchically over seasons.
@@ 3.2 Susceptible-Infectious-Recovered (SIR) compartmental model
The SIR compartmental model is a mathematical model used for modeling disease outbreaks and was introduced by Kermack and McKendrick Kermack and McKendrick (1927). Since then, compartmental models have become standard for modeling infectious diseases Allen et al. (2008), and many extensions have been made and studied (Simon, 2020; Allen, 2017; Van den Driessche, 2008, for example). The SIR mathematical model includes three compartments and assumes that at any time $t>0$ every individual in a closed population belongs to exactly one compartment. The three compartments are susceptible ($S$), infectious ($I$), and recovered ($R$), and their interaction over the course of an outbreak is described by the differential equations in (2).
@@
Here $S$, $I$, and $R$ represent the proportion of the population in each compartment such that $S+I+R=1$ for all $t$. The trajectory is determined by the disease transmission rate $\beta>0$ and the recovery rate $\delta>0$. Respectively, these may be thought of as the expected proportion of susceptible individuals who will be infected by an infectious individual and the expected rate of recovery to an immune state for a newly infected person. Whether or not a disease outbreak is classified as an epidemic is determined by the initial susceptible population $S_{0}$, or the susceptible population at time 00, and the parameter $\rho=\delta/\beta$. If $S_{0}/\rho>1$, the outbreak is considered an epidemic. It is non-epidemic if $S_{0}/\rho\leq 1$ Osthus et al. (2019). Figure 2 shows the trajectory of the three compartments of an SIR model where $S_{0}$ and $\rho$ were selected to match an outbreak that is an epidemic. In the case where $S_{0}\leq\rho$, the trajectory for the $I$ compartment would never be increasing. The increase to a peak and subsequent decrease in the $I$ compartment of figure 2 suggest it is reasonable to model ILI by this compartment. Thus in modeling the ILI data, we consider the data to be analogous to the $I$ proportion of the population.
@@ 3.3 Asymmetric Gaussian (ASG) function
The ASG function is another example of a nonlinear function which can approximate the trajectory of the flu outbreak. The ASG was previously used by Ulloa to model and forecast ILI Ulloa (2019), and it has been used to model vegetation growth and satellite sensor data Lewis-Beck et al. (2020); Jonsson and Eklundh (2002); Hird and McDermid (2009); Beck et al. (2006); Atkinson et al. (2012). The ASG is a modification of the asymmetric Gaussian distribution Wallis (2014) and is characterized by its rise to a peak and fall from that peak which may not occur at the same rate, as shown in figure 3. The ASG function is denoted as $ASG_{\theta}(w)$ where $\theta=(\lambda,\nu,\mu,\sigma_{1}^{2},\sigma_{2}^{2})$, $\nu>0$, $\lambda>0$, $\mu\in(-\infty,\infty)$, $\sigma_{1},\sigma_{2}>0$ and $w\in(1,...,W)$ is week. The function is defined in (3).
@@
For modeling in this manuscript, we use a slightly reparameterized version of the function in (4), where $\eta=\nu-\lambda>0$. This constraint guarantees that the function has a peak greater than $\lambda$.
@@ 3.4 Model discrepancy
The SIR and ASG functions are useful for capturing the main trend of the ILI data, but as Osthsu et al. Osthus et al. (2019) point out there may be systematic behavior in the ILI that these or other possible functions may not capture. As noted in section 2, figure 1 shows a regular peak at week 22 of the flu season. Figures 4 and 5 are used together to illustrate the systematic discrepancy from a fitted function. Figure 4 shows the US ILI percentage for all flu seasons from 2010 to 2022 excluding 2020 with a best fit ASG function plotted over the ILI. The fits for each season were made by obtaining the maximum likelihood estimate (MLE) of a model given the ILI data where we assume the ASG function is the mean parameter of a Beta distributed random variable. Figure 5 shows the discrepancy between the model fit and the data for the same seasons. The grey lines show the difference between the data and the functions from figure 4 for each season, and the black line is the average by week over all seasons. The lines show that the ASG function typically underpredicts week 22 and overpredicts week 23. Perhaps for other weeks, like week 30 for example, there also tends to be systematic behavior not captured by the ASG function.
@@
The term $\gamma_{w}$, where $w$ is the season week, is included in (1) to capture the per week discrepancy between ILI and the function. Modeling discrepancy has been used in uncertainty analysis of simulators to capture systematic differences between mathematical models and reality Ma et al. (2022); Brynjarsdóttir and O’Hagan (2014); Arendt et al. (2012); Kennedy and O’Hagan (2001). Modeling discrepancy can lead to overfitting, particularly in forecasting scenarios, and may also lead to identifiability issues. Thus, care must be taken in setting parameter constraints as well as in the selection of prior distributions Osthus et al. (2019); Brynjarsdóttir and O’Hagan (2014). Modeling of the discrepancy for ILI was done by Osthus et al. during the 2015 and 2016 flu seasons where their model outperformed all others in the CDC flu forecasting challenge Osthus et al. (2019). As in Osthus’s model, $\gamma_{w}$ is modeled as a reverse random walk, as shown in (5).
$\gamma_{w}|\gamma_{w+1}\overset{ind}{\sim}N(\gamma_{t+1},\sigma^{2}_{\gamma}),\quad\gamma_{W}\sim N(0,\sigma^{2}_{\gamma_{W}})$ (5)
The idea for using the reverse random walk is that there are several previous seasons of ILI data, and assuming the random walk captures systematic behavior, fitting it hierarchically over seasons can assist in predicting future behavior in the current season. Reverse random walks have also been used with success in election forecasting and other flu forecasting models Osthus and Moran (2021); Osthus et al. (2019); Linzer (2013). The sum to zero constraint $-\gamma_{1}=\sum_{w=2}^{W}\gamma_{w}$ is imposed on (5) to improve identifiability.
3.5 ILI forecasts
Model (1) is fit via Bayesian posterior updating. Future ILI forecasts are obtained via the posterior predictive distribution where for week $w$, the predictive distribution is obtained by integrating over the parameters $\pi_{s}$, $\kappa_{s}$, $\sigma^{2}_{\gamma}$, and $\sigma^{2}_{\gamma_{W}}$ as in (6) where $p(\bm{\pi}_{s},\kappa_{s},\sigma^{2}_{\gamma},\sigma^{2}_{\gamma_{W}}|\textbf{ILI})$ is the density function of the posterior distribution for the model parameters. If the current week is $w^{*}$ then the desired forecasts are for weeks $w^{*}+i$ where $i$ is a positive integer.
[…]
@@ 3.6 Hospitalization model
The second component for forecast modeling is based on the relationship between hospitalizations and ILI and is defined in (7). This is an example of an autoregressive model with exogenous variables where the autoregressive lag is one (ARX(1)) Raftery et al. (2010); Ljung (1987). Figure 6 shows scatterplots of the difference between hospitalizations and scaled 1 week lags by ILI percentage at the US national level. The points in the left plot are from hospitalization differences whereas in the right the hospitalizations were log transformed. The positive and slightly curved relationship suggests model (7) is a reasonable choice.
Figure 6: Scatterplots of national hospitalizations (left) or log hospitalizations (right) lag differences by ILI % where the lag is scaled by $\phi$. Points are colored by season. For hospitalizations $\phi=0.6$, and for log hospitalizations $\phi=0.77$.
@@
In model (7), $H_{s,w}$ is the number of hospitalizations for week $w$ in season $s$, $\epsilon_{s,w}$ is an error term distributed according to some distribution $D_{s}$ with mean parameter 0, scale parameter $\sigma_{\epsilon_{s}}$, and the additional parameter $\omega_{s}$ is the degrees of freedom parameter when $D_{s}$ belongs to the location-scale t (LST) family.
This model is for any location, and for fitting purposes $ILI_{s,w}$ is always multiplied by $P$ which is proportional to the population of the state or territory, in this case the total population divided by 50,000. This is done as a means of scaling so that the prior distribution assigned to $\bm{\alpha}_{s}=(\alpha_{0s},\alpha_{1s},\alpha_{2s})$, $\sigma_{\epsilon_{s}}$, and $\omega_{s}$ might reasonably be the same for all states.
Like the ILI model, the hospitalization model in (7) is also fit via Bayesian posterior updating. To obtain forecasts for $H_{s,w^{*}+i}$, the ILI posterior predictive distribution is used along with the posterior distribution for the parameters in (7). We considered three scenarios for model (7). A model where $D_{s}$ belongs to the normal family (NORM), $D_{s}$ belongs to the LST family, and one where $H_{s,w}$ is replaced with $\text{log}(H_{s,w}+1)$ and $D_{s}$ is from a normal family, or $H_{s,w}+1$ is lognormally distributed (LNORM). The population value $P$ is excluded from the LNORM model, and we set $\bm{\alpha}_{22}=\bm{\alpha}_{23}$ to help with fitting. In the LNORM model, if the linear parameters are not set the same for seasons 22 and 23, the final variance was more prone to be extreme. Besides varying the distribution family of hospitalizations, we also considered $\alpha_{2s}=0$ or there is no quadratic ILI term.
@@ 3.7 Prior selection
The priors selected for the ILI data model under both the SIR and ASG models largely follow the prior selections in Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) with a few exceptions where changes improved numerical stability and/or we felt the adjusted prior made more sense for the problem. For model (1), parameters which are common even when using different functions of $f_{\theta_{s}}(w)$ are $\kappa_{s}$, $\sigma_{\gamma}^{2}$, and $\sigma_{\gamma_{W}}^{2}$. For the SIR function $\theta_{s}=(S_{0s},I_{0s},R_{0s},\alpha_{s},\rho_{s})$, and for the ASG function $\theta_{s}=(\alpha_{s},\eta_{s},\mu_{s},\sigma_{1s}^{2},\sigma_{2s}^{2})$. For the hospitalization model in (7) the parameter to be estimated is $\Psi=(\alpha_{0s},\alpha_{1s},\alpha_{2s},\phi,\sigma_{\epsilon_{s}},\omega_{s})$.
@@
The priors assigned were mostly noninformative, though in certain cases the prior distributions were selected for numerical stability as was the case for $\sigma_{\gamma}^{2}$ and $\sigma_{\gamma_{W}}^{2}$. For these two scale parameters only, rather than assigning a half-normal prior to the standard deviation parameters, as recommended by Gelman Gelman (2006), the priors were assigned to the variance parameters. Univariate parameters were assigned either a normal distribution prior if the support is on $\mathbb{R}$, a half-normal prior if the support is nonnegative, or a truncated-normal prior to match a more specific support. Under the ASG model, $\theta_{s}$ is modeled hierarchically over seasons so that for each season the transformed parameter $T(\theta_{s})\sim N(\theta,\Sigma$) and priors distributions are assigned to $\theta$ and $\Sigma$.
Additional prior constraints were made to improve parameter identifiability. In Osthus et al. Osthus et al. (2019) the initial value of the susceptible population compartment of the SIR model was set to $S_{0}=0.9$. The parameters $I_{0s}$, $\beta_{s}$, and $\rho_{s}$ were assigned informative priors. To improve identifiability when $f_{\theta_{s}}(w)=ASG_{\theta_{s}}(w)$ in (1) we followed a modular Bayesian approach for fitting the parameters. A modular Bayesian approach involves multiple steps of parameter fitting where some parameters may be estimated without priors via maximum likelihood estimation or other means. Fitting the rest of the model parameters involves assigning priors and conditioning on the previously fit parameters. This has been done in computer modeling to improve identifiability and other issues, though Liu et al. Liu et al. (2009) warn this approach is not probabilistically sound if parameter inference is a priority Jiang et al. (2015); Arendt et al. (2012); Liu et al. (2009). We carried out the modular fit by first estimating the maximum likelihood estimate (MLE) for the parameter $\lambda_{s}$ for each season in (4) and plugging in the MLE as a fixed value.

### v2

@@ 3.1 ILI Model
The proposed model for ILI for any location is given in (1). Here $ILI_{s,w}$ is the ILI for flu season $s$ and week w=1,2,,Ww=1,2,...,W, where $W=52$ or $W=53$, depending on how many Sundays there are in a given season. The ILI is a proportion, so the Beta random variable is a natural selection for modeling. Under the parameterization of the Beta distribution used in (1) the expected value is $\pi_{s,w}$ and the variance is $\pi_{s,w}(1-\pi_{s,w})/(1+\kappa_{s})$, making $\kappa_{s}$ a scale parameter. The nonlinear function $f_{\theta_{s}}(w)$ captures the trajectory of the ILI, and $\gamma_{w}$ is a discrepancy term shared by all seasons which captures systematic patterns not captured by $f_{\theta_{s}}(w)$. The additional discrepancy term $\upsilon_{s,w}$ captures additional season specific patterns.
@@
$\text{logit}(\pi_{s,w})$ $=f_{\theta_{s}}(w)+\gamma_{w}+\upsilon_{s,w}$
In Osthus et al. Osthus et al. (2019) $f_{\theta_{s}}(w)=\text{logit}(I_{s,w})$ where $I_{s,w}$ is the infectious compartment of the SIR model from (2) in section 3.2. In Ulloa Ulloa (2019), $f_{\theta_{s}}(w)=ASG_{\theta}(w)$ from (3) in section 3.3. In Ulloa Ulloa (2019), modeling is done hierarchically over seasons. Herein we also adopt the hierarchical modeling so as to borrow information contained in the data across all seasons.
Model (1) was fit independently for each location. This decision was made largely due to the limited time given by FluSight to produce forecasts. That is, during the competition, data were typically released on Wednesday afternoons and forecasts were due by the evening of the same day giving only a few hours to produce forecasts. A natural extension of the model would be to allow for states to borrow information from neighboring states.
@@ 3.2 Susceptible-Infectious-Recovered (SIR) compartmental model
The SIR compartmental model is a mathematical model used for modeling disease outbreaks and was introduced in 1927 by Kermack and McKendrick Kermack and McKendrick (1927). Since then, compartmental models have become standard in modeling infectious diseases Allen et al. (2008), and many extensions have been made and studied (Simon, 2020; Allen, 2017; Van den Driessche, 2008, for example). The SIR mathematical model includes three compartments and assumes that at any time $t>0$ every individual in a closed population belongs to exactly one compartment. The three compartments are susceptible ($S$), infectious ($I$), and recovered ($R$), and their interaction over the course of an outbreak is described by the differential equations in (2).
@@
Here $S$, II, and $R$ represent the proportion of the population in each compartment such that $S+I+R=1$ for all $t$. The trajectory is determined by the disease transmission rate $\beta>0$ and the recovery rate $\delta>0$. Respectively, these may be thought of as the expected proportion of susceptible individuals who will be infected by an infectious individual and the expected rate of recovery to an immune state for a newly infected individual. Whether or not a disease outbreak is classified as an epidemic is determined by the initial susceptible population $S_{0}$, or the susceptible population at time 0, and the parameter $\rho=\delta/\beta$. If $S_{0}/\rho>1$, the outbreak is considered an epidemic. It is non-epidemic if $S_{0}/\rho\leq 1$ Osthus et al. (2019). Figure 2 shows the trajectory of the three compartments of an SIR model where $S_{0}$ and $\rho$ were selected to match an outbreak that is an epidemic. In the case where $S_{0}\leq\rho$, the trajectory for the $I$ compartment will be to never increase. The increase to a peak and subsequent decrease in the $I$ compartment of figure 2 illustrate why it is reasonable to use the SIR model to partially capture ILI, although it is unlikely to completely capture the true data generating mechanism. Thus the use of modeling discrepancy.
@@ 3.3 Asymmetric Gaussian (ASG) function
The ASG function is another example of a nonlinear function which can approximate the trajectory of the flu outbreak. The ASG was previously used by Ulloa to model and forecast ILI Ulloa (2019). It has also been used to model vegetation growth and satellite sensor data Lewis-Beck et al. (2020); Jonsson and Eklundh (2002); Hird and McDermid (2009); Beck et al. (2006); Atkinson et al. (2012). The ASG is a modification of the asymmetric Gaussian distribution Wallis (2014) and is characterized by its rise to a peak and fall from that peak, which rise and fall may not occur at the same rate, as shown in figure 3. The ASG function is denoted as $ASG_{\theta}(w)$ where $\theta=(\lambda,\nu,\mu,\sigma_{1}^{2},\sigma_{2}^{2})$, $\nu>0$, $\lambda>0$, $\mu\in(-\infty,\infty)$, $\sigma_{1},\sigma_{2}>0$ and $w\in(1,...,W)$ is week. The function is defined in (3).
@@
In this manuscript, we use a slightly reparameterized version of function (3) shown in (4), where $\eta=\nu-\lambda>0$. This constraint guarantees that the function has a peak greater than $\lambda$.
@@ 3.4 Model discrepancy
The SIR and ASG functions are useful for capturing the main trend of the ILI data, but as Osthsu et al. Osthus et al. (2019) point out there may be systematic behavior in the ILI that these or other possible functions may not capture. Figures 4 and 5 are used together to illustrate the systematic discrepancy from a fitted function to ILI data. Figure 4 shows the US ILI percentage for all flu seasons from 2010 to 2022 excluding 2020 with a best fit ASG function plotted over the ILI. The fits for each season were made by computing the maximum likelihood estimate (MLE) of a model given the ILI data where we assume the ASG function is the mean parameter of a Beta distributed random variable. Figure 5 shows the discrepancy between the model fit and the data for the same seasons. The grey lines show the difference between the data and the functions from figure 4 for each season, and the black line is the average discrepancy over all seasons. The lines show that the ASG function typically underpredicts week 22 and overpredicts week 23. Perhaps for other weeks, week 30 for example, there tends to be additional systematic behavior not captured by the ASG function.
@@
Modeling discrepancy has been used in uncertainty analysis of simulators to capture systematic differences between mathematical models and reality Ma et al. (2022); Brynjarsdóttir and O’Hagan (2014); Arendt et al. (2012); Kennedy and O’Hagan (2001). Modeling discrepancy can lead to overfitting, particularly in forecasting scenarios, and may also lead to identifiability issues. Thus, care must be taken in setting parameter constraints as well as in the selection of prior distributions Osthus et al. (2019); Brynjarsdóttir and O’Hagan (2014). For instance, when setting $f_{\theta_{s}}(w)$ to be the ASG function in model (1), we found the parameter $\lambda$ from (4) difficult to estimate precisely because of lack of identifiability. This was resolved by assigning to $\lambda$ a more informative prior distribution. Modeling of the discrepancy for ILI was done by Osthus et al. during the 2015 and 2016 flu seasons where their model outperformed all others in the CDC flu forecasting challenge Osthus et al. (2019). The terms $\gamma_{w}$ and $\upsilon_{s,w}$, where $s$ is the season and $w$ is the season week, are included in (1) to capture the per week discrepancy between ILI and the function. Here $\gamma_{w}$ is shared across all seasons, whereas $\upsilon_{s,w}$ is specific to season $s$. To assist in estimating $\gamma_{w}$, $\upsilon_{s,w}$ is only included for the current season being forecast and none of the previous seasons. As in Osthus’s model, $\gamma_{w}$ and $\upsilon_{s,w}$ are modeled as reverse random walks, as shown in (5).
$\gamma_{w}|\gamma_{w+1}$ $\overset{ind}{\sim}N(\gamma_{w+1},\sigma^{2}_{\gamma}),\quad\gamma_{W}\sim N(0,\sigma^{2}_{\gamma_{W}})$ (5)
$\upsilon_{s,w}|\upsilon_{s,w+1}$ $\overset{ind}{\sim}N(\upsilon_{s,w+1},\sigma^{2}_{\upsilon})$
The idea for using the reverse random walk is that there are several previous seasons of ILI data, and assuming the random walk captures systematic behavior, fitting it hierarchically over seasons can assist in predicting future behavior in the current season. Reverse random walks have also been used with success in election forecasting and other flu forecasting models Osthus and Moran (2021); Osthus et al. (2019); Linzer (2013). The sum to zero constraint $\gamma_{1}=\sum_{w=2}^{W}\gamma_{w}$ is imposed on (5) to improve identifiability as well as the constraint $\upsilon_{s,W}=0$.
@@ 3.5 Hospitalization model
The second component for forecast modeling is based on the relationship between hospitalizations and ILI and is defined in (6). This is an example of an autoregressive model with exogenous variables where the autoregressive lag is one (ARX(1)) Raftery et al. (2010); Ljung (1987).
@@
Figure 6 shows US national level data which helps illustrate why this modeling decision was made. The figure shows scatterplots of ILI percentage and hospitalizations on the top row. The bottom row shows scatterplots of the difference between hospitalizations and 1 week lags by ILI percentage. The plots in the right column are of the logarithm of hospitalizations. The lines through the points are drawn by connecting hospitalizations for consecutive weeks during the flu season. These plots show that a linear or quadratic relationship between ILI percentage and hospitalizations does not capture much of the temporal correlation in hospitalizations as shown by the obvious loops drawn in the plot. However, when including a single autoregressive lag of hospitalizations in the model, much of the time correlation is accounted for as shown by the tighter and more frequent crossing of lines in the bottom row. We found that models with more than 1 autoregressive lag showed little or no forecast improvement over 1 lag.
Figure 6: Scatterplots of national hospitalizations (left) or log hospitalizations (right) lag differences by ILI % where the lag is scaled by $\phi$. Points are colored by season. For hospitalizations $\phi=0.6$, and for log hospitalizations $\phi=0.77$. Lines through the points follow the weekly progression of influenza.
In model (6), $H_{s,w}$ is the number of hospitalizations for week $w$ in season $s$, $\epsilon_{s,w}$ is an error term distributed according to some distribution $D_{s}$ with mean parameter 0, scale parameter $\sigma_{\epsilon_{s}}$, and the additional parameter $\omega_{s}$ which may be used as degrees of freedom when $D_{s}$ belongs to the location-scale t (LST) family.
Like the ILI model (1), the hospitalization model was fit independently for each location. For fitting purposes $ILI_{s,w}$ is always multiplied by $P$ which is proportional to the population of the state or territory, in this case the total population divided by 50,000. This is done as a means of scaling so that the prior distribution assigned to $\bm{\alpha}_{s}=(\alpha_{0s},\alpha_{1s},\alpha_{2s})$, $\sigma_{\epsilon_{s}}$, and $\omega_{s}$ might reasonably be the same for all locations.
Like the ILI model, the hospitalization model in (6) is also fit via Bayesian posterior updating. To obtain forecasts for $H_{s,w^{*}+i}$, the ILI posterior predictive distribution is used along with the posterior distribution for the parameters in (6). We let $D_{s}$ belong to the normal family, though in the supplementary materials we present a real forecast analysis where we considered other distribution families including the LST and lognormal families.
@@ 3.6 Prior selection
The prior distributions selected for the ILI data model under both the SIR and ASG models largely follow the prior distribution selections in Osthus et al. Osthus et al. (2019) and Ulloa Ulloa (2019) with a few exceptions where selecting different priors improved identifiability, numerical stability, and/or we felt the adjusted prior made more sense for the problem. For model (1), parameters which are common even when using different functions of $f_{\theta_{s}}(w)$ are $\kappa_{s}$, $\sigma_{\gamma}^{2}$, $\sigma_{\gamma_{W}}^{2}$, and $\sigma_{\upsilon}$. For the SIR function $\theta_{s}=(S_{0s},I_{0s},R_{0s},\alpha_{s},\rho_{s})$, and for the ASG function $\theta_{s}=(\lambda_{s},\eta_{s},\mu_{s},\sigma_{1s}^{2},\sigma_{2s}^{2})$. For the hospitalization model in (6) the parameter to be estimated is $\Psi=(\alpha_{0s},\alpha_{1s},\alpha_{2s},\phi,\sigma_{\epsilon_{s}},\omega_{s})$.
@@
The priors assigned were mostly noninformative, though in certain cases the prior distributions were selected for numerical stability as was the case for $\sigma_{\gamma}^{2}$ and $\sigma_{\gamma_{W}}^{2}$. For these two scale parameters only, rather than assigning a half-normal prior to the standard deviation parameters, as recommended by Gelman Gelman (2006), the priors were assigned to the variance parameters. Univariate parameters were assigned either a normal distribution prior if the support was on $\mathbb{R}$, a half-normal prior if the support is nonnegative, or a truncated-normal prior to match a more specific support. Under the ASG model, $\theta_{s}$ was modeled hierarchically over seasons so that for each season the transformed parameter $T(\theta_{s})\sim N(\theta,\Sigma$) and priors distributions are assigned to $\theta$ and $\Sigma$ where $T(\theta_{s})=(\lambda_{s},\text{log}(\eta_{s}),\mu_{s},\text{log}(\sigma_{1s}^{2}),\text{log}(\sigma_{2s}^{2}))$.
Additional prior constraints were made to improve parameter identifiability. Just as was done in Osthus et al. Osthus et al. (2019), we set the initial value of the susceptible population compartment of the SIR model was set to $S_{0}=0.9$. The parameters $I_{0s}$, $\beta_{s}$, and $\rho_{s}$ were assigned informative priors. In a previous version of this manuscript, when fitting an ASG model with discrepancy components, rather than estimating all parameters through posterior updating, identifiability was encouraged by setting the parameter $\lambda_{s}$ from (4) to be equal to its MLE. In this updated manuscript, however, we use empirical Bayes and center the first component of the hierarichal prior $\theta$ to be the mean of MLEs for $\lambda_{s}$ across all seasons and set the variance parameter so the prior was tight around the mean. This led to improved mixing of posterior draws and better forecasts. The specific prior distributions assigned to each parameter are listed in the supplementary materials.
3.7 Hospitalization Forecasts Using the Bayesian Cut
Rather than combining the ILI model in equation (1) with the hospitalization model in equation (6) and fitting a single posterior distribution, both models were fit separately and 1 to 4 week ahead ILI forecasts from model (1) were combined with the posterior predictive distribution of model (6) to obtain 1 to 4 week ahead hospitalization forecasts. In this scheme, the posterior distribution of model (1) parameters was not directly influenced by the hospitalization data and the posterior distribution of model (6) parameters was not directly influenced by model (1) parameters. Hence there was limited feedback between the two components, but propagation of uncertainty in ILI forecasts was still permitted in the estimating of hospitalization forecasts.
This two-component modeling is remeniscent of the Bayesian cut described in Plummer (2015) and Nott et al. (2023). Cutting a model into two component models or “modules“ may be justfied for several reasons including to reduce time to fit the model, avoiding mixing issues of posterior sampling distributions, improving predictive performance or to prevent model misspecification in one model to influence the other Nott et al. (2023); Jacob et al. (2020, 2017); Plummer (2015).
The decision to cut the forecast model into two component models was made for several reasons. When fitting the fully Bayesian joint model, the time to fit a single model using Markov chain Monte Carlo (MCMC) sampling and obtain forecasts was sometimes many hours whereas fitting the two models separately took a fraction of the time. When assessing convergence of model parameters, it was rare that posterior sampling chains for the joint model showed convergence even after 55,000 iterations, whereas the component model fits gave no signs which raised concern for convergence. And most importantly, the forecasts from the component models generally outperformed those of the joint model in terms of minimizing proper scoring rules.
Model (1) was fit via Bayesian posterior updating. Future ILI forecasts were obtained via the posterior predictive distribution where for week $w$, the predictive distribution was obtained by integrating over the parameters $\pi_{s}$, $\kappa_{s}$, $\sigma^{2}_{\gamma}$, and $\sigma^{2}_{\gamma_{W}}$ as in (7) where $p(\bm{\pi}_{s},\kappa_{s},\sigma^{2}_{\gamma},\sigma^{2}_{\gamma_{W}}|\textbf{ILI})$ is the density function of the posterior distribution for the model parameters. If the current week was $w$ then the desired forecasts were for weeks $w+i$ where $i$ is a positive integer. Here $\widetilde{ILI}_{s,w+i}$ represents the predicted or forecasted ILI at $i$ weeks in the future.
$p(\widetilde{ILI}_{s,w+i}|\textbf{ILI})=\int\int\int\int p(\widetilde{ILI}_{s,w+i}|\bm{\pi}_{s},\kappa_{s},\sigma^{2}_{\gamma},\sigma^{2}_{\gamma_{W}})p(\bm{\pi}_{s},\kappa_{s},\sigma^{2}_{\gamma},\sigma^{2}_{\gamma_{W}}|\textbf{ILI})d\bm{\pi}_{s}d\kappa_{s}d\sigma^{2}_{\gamma}d\sigma^{2}_{\gamma_{W}}$ (7)
Hospitalization forecasts $\widetilde{H}_{s,w+i}$ at week $w+i$ were then obtained by combining the posterior predictive distribution of model (6) with (7) and integrating over $\widetilde{ILI}_{s,w+i}$ as in (8). Here, $\bm{\alpha}$ represents all regression and variance parameters in model (6).
$p(\tilde{H}_{s,w+i}|\textbf{ILI},\textbf{H},\bm{\alpha})$ $=\int p(\tilde{H}_{s,w+i}|\textbf{H},\textbf{ILI},\widetilde{ILI}_{s,w+i},\bm{\alpha})p(\widetilde{ILI}_{s,w+i}|\textbf{ILI})d\widetilde{ILI}_{s,w+i}$ (8)